OpenStudy (anonymous):

Solve for x 2 - 4/x^2 = -2/x

OpenStudy (shamim):

2x^2-4=-2x

OpenStudy (shamim):

2x^2+2x-4=0

OpenStudy (shamim):

2x^2+4x-2x-4=0

OpenStudy (shamim):

R u able to do the rest

OpenStudy (anonymous):

@rathanreddy

OpenStudy (anonymous):

\[x^{3}+x ^{2}-x=0 this is the equation comes when we solve \it now if you have any options then substitute \in this equation if the answer is 0 then that option is correct\]

OpenStudy (anonymous):

sorry? there are no spaces

OpenStudy (anonymous):

can you read it

OpenStudy (anonymous):

should i type it again

OpenStudy (anonymous):

no its okay I got it

OpenStudy (anonymous):

k

OpenStudy (anonymous):

there will be 3 answers one is 0 and other's you find

OpenStudy (anonymous):

x=-2 or 1 @Alexis2332

OpenStudy (anonymous):

im not sure how to work it out though

OpenStudy (anonymous):

do you have any options or i should give the answers

OpenStudy (nincompoop):

try to give the answer and let me see you banned :)

OpenStudy (anonymous):

 @Alexis2332